# Semantically-Shifted Incremental Adapter-Tuning is A Continual ViTransformer

### v1

@@ 3.3 Semantic shift estimation without past samples
Due to the selective updating of classifiers corresponding to the current task during training, the classifiers across different learning sessions are not fully aligned in the same feature space. To further optimize classifiers, we store the prototypes after training the backbone and local classifier. However, as the backbone is trained incrementally with new classes, the feature distribution of old classes undergoes changes. Retraining the classifier with the previous prototypes is sub-optimal. Since the feature representability of the backbone updates over time, using outdated features may not effectively retrain a unified classifier. To solve this problem, we update the feature distribution of old classes by computing the semantic shift over the learning process. We follow SDC [42] to estimate the semantic shift of old prototypes without access to past samples.
@@
While we do not have access to data from the old class $c$, we can only estimate the shift of current task categories on old and new models. The semantic shift of current samples between two sessions can be represented as
$\delta_{i}^{t-1\to t}=e_{i}^{t}-e_{i}^{t-1},$ (8)
where $e$ denotes the embedding of one sample in the current task $t$. We can compute $e_{i}^{t-1}$ at the start of the current task with the model trained in task $t-1$. After training on the new task, we compute $\delta_{i}^{t-1\to t}$ and use it to estimate $\Delta_{c}^{t-1\rightarrow t}$. We compute the shift as
$\widetilde{\Delta}_{c}^{t-1\to t}=$ $\frac{\sum\alpha_{i}\delta_{i}^{t-1\to t}}{\sum\alpha_{i}},c\notin C^{t},$ (9)
$\alpha_{i}=$ $\mathbf{e}^{-\frac{||e_{i}^{t-1}-\varphi_{c}^{t-1}||}{2\sigma^{2}}},$
where $\sigma$ is the standard deviation of the distribution of class $c$; $C^{t}$ denotes classes learned in the current session. Before retraining the classifier, we update the prototypes with
$\left\{\begin{array}[]{lc}\varphi_{c}=\varphi_{c}^{t-1}+\widetilde{\Delta}_{c}^{t-1\to t}&,c\notin C^{t}\\ \varphi_{c}=\frac{1}{N_{c}}\sum_{i}e_{c}&,c\in C^{t},\\ \end{array}\right.$ (10)
where $N_{c}$ denotes the number of images in class $c$.
@@ 3.4 Unified classifier training
$\mathcal{L}(\theta_{cls},\mathcal{V}_{c})=-\sum_{i=1}^{S_{n}*C}log\frac{\mathbf{e}^{(\theta_{cls}^{j}(v_{i}))}}{\sum_{k\in C}\mathbf{e}^{(\theta_{cls}^{k}(v_{i}))}},\vspace{-1mm}$ (11)
@@
Method Params Split-ImageNetR Split-ImageNetA CUB200 CIFAR100
@@
SSIAT (Ours) 1.19M $79.38_{\pm 0.59}$ $83.63_{\pm 0.43}$ $62.43_{\pm 1.63}$ $70.83_{\pm 1.63}$ $88.75_{\pm 0.38}$ $93.00_{\pm 0.90}$ $91.35_{\pm 0.26}$ $94.35_{\pm 0.60}$
Table 1: Experimental results on four CIL benchmarks. All other methods are reproduced using the same seeds for a fair comparison.
Figure 5: The performance of each learning session on four datasets. (a) ImageNetR; (b) ImageNetA; (c) CUB200; (d) CIFAR100. These curves are plotted by calculating the average performance across three different seeds for each incremental session.
@@ 4.1 Datasets and Evaluation Protocols
Dataset: We evaluate our method on four commonly-used CIL benchmarks and one cross-domain CIL dataset. We randomly split the dataset into 10 or 20 learning tasks. CIFAR100 [18] is a widely used dataset in CIL which consists of 60000 images, belonging to 100 different categories. CUB200 [33] is a dataset that contains approximately 11,788 images of 200 bird species with fine-grained class labels. Additionally, we also follow recent work [50, 45] to use the other three datasets which have a large domain gap with pre-training data. ImageNetR [10] consists of 30,000 images with 200 categories. Although its categories overlap with ImageNet-21K [29], the images belong to a different domain. ImageNetA [11] is a real-world dataset that consists of 200 categories. This dataset exhibits significant class imbalance, with some categories having only a few training samples. VTAB [44] is a complex dataset that consists of 19 tasks covering a broad spectrum of domains and semantics. We follow previous work [50] to select 5 tasks to construct a cross-domain CIL dataset.
@@
SSIAT (Ours) $75.67_{\pm 0.14}$ $82.30_{\pm 0.36}$ $59.16_{\pm 1.03}$ $68.45_{\pm 1.92}$
Table 2: Experimental results for long-sequences (20 incremental sessions) on ImageNetR and ImageNetA dataset.
Implementation details: We use ViT-B/16 [5] as the pre-trained model, which is pre-trained on ImageNet-21K [29]. The initial learning rate is set as 0.01 and we use the cosine Anneal scheduler. In our experiments, we train the first session for 20 epochs and 10 epochs for later sessions. Following previous papers [50, 45], we use common evaluation metrics in CIL. Specifically, we report the last session accuracy $\mathcal{A}_{Last}$ and average accuracy of the whole incremental sessions $\mathcal{A}_{Avg}=\frac{1}{T}\sum_{i=1}^{T}\mathcal{A}_{i}$. We utilize three different seeds to generate three different class orders for evaluating various methods. We report the mean and standard deviation based on the three experiments. See codes [^1] https://github.com/HAIV-Lab/SSIAT.
@@ 4.2 Experiment Results
For a fair comparison, we compare our methods with SOTA CIL methods based on the pre-trained vision transformer model. We compare our methods with prompt-based methods L2P [52], DualPrompt [37], CODAPrompt [31], fine-tuning methods SLCA [45], and adapter-based method [50, 8, 6]. Tab. 1 shows $\mathcal{A}_{Avg}$ and $\mathcal{A}_{Last}$ with three different seeds on four CIL benchmarks.
@@
SSIAT (Ours) 96.10 92.71 94.09 93.68 94.50 94.21
Table 3: Experimental results for different methods on VTAB dataset which contain 5 datasets from different domains.
@@
Table 4: Experimental results for baselines with different efficient tuning methods on CIFAR100. We report the overall performance of each session and the average performance.
ImageNetR &\&& ImageNetA: We report the performance on ImageNetR and ImageNetA in Tab. 1. These two datasets are more difficult due to the domain gap with the pre-training data. It can be seen that the performance of each method on these two datasets is lower than CIFAR100 and CUB200. Besides, we can see that SLCA outperforms other previous methods significantly on these two datasets. Notably, SLCA achieves an impressive last accuracy on ImageNetR, surpassing the other methods. In contrast, our method achieves SOTA-level performance on both datasets with fewer tuning parameters. Based on Fig. 5, the performance of our method is slightly higher than SLCA in several learning sessions with fewer tuning parameters on the ImageNetR dataset. On the ImageNetA dataset, our method achieves the last accuracy of 62.43%, surpassing SLCA by 1.39%. The average accuracy across all sessions is 70.83%, showing a 2% improvement.
Additionally, we evaluate the performance of each method under the condition of long sequences. In this setting, each session consists of only 10 classes, and the results are summarized in Tab. 2. Our method also maintains excellent performance in terms of $\mathcal{A}_{Last}$ and $\mathcal{A}_{Avg}$. The performance of SLCA is highly dependent on the class order in which the training data appears, resulting in a substantial variance in $\mathcal{A}_{Last}$ on ImageNetA. In contrast, the Adam-based methods remain relatively stable in long-sequence settings. For Adam-SSF, the long sequence only leads to a nearly $2\%$ performance drop in ImageNetR. However, for SLCA, its performance drops by $5\%$ on ImageNetR and nearly $10\%$ on ImageNetA. In comparison, our method demonstrates excellent stability on long sequences and outperforms other methods by a large margin.
@@
w/ SSCA 82.66 80.60 75.91 72.41 71.56 69.01 66.10 64.60 63.00 62.43 70.83
Table 5: Ablation results for unified classifier training and semantic shift estimation on ImageNetA. We report the overall performance of each session and the average performance. We run the experiments with three seeds and reported the average performance.
Method CIFAR ImageNetR ImageNetA CUB
@@
Table 6: Linear probing results of different training ways on four datasets. We retrain the classifier using all the data on the fixed-trained backbone.
@@
Table 7: Experimental results of different adapter structures. We report the average performance and standard deviation.
VTAB: VTAB is a cross-domain CIL dataset where each task provides training data from a different domain. Based on the results presented in Tab. 3, it can be observed that both SLCA and our method perform well in cross-domain CIL. Specifically, in the last incremental stage, our method achieves an accuracy that is $12\%$ higher than the Adam-based methods. Adam-based methods only perform fine-tuning in the first task and are not able to adapt well to subsequent tasks on the cross-domain dataset.

### v2

@@ 4.2 Semantic shift estimation without past samples
Due to the selective updating of classifiers corresponding to the current task during training, the classifiers across different learning sessions are not fully aligned in the same feature space. To further optimize classifiers, we store the prototypes after training the backbone and local classifier. However, as the backbone is trained incrementally with new classes, the feature distribution of old classes changes. Retraining the classifier with the previous prototypes is sub-optimal. Since the feature representability of the backbone updates over time, using outdated features may not effectively retrain a unified classifier. To solve this problem, we update the feature distribution of old classes by computing the semantic shift over the learning process. Inspired by the previous work SDC [33], we estimate the semantic shift of old prototypes without access to past samples.
@@
While we do not have access to data from the old class $c$, we can only estimate the shift of current task categories on old and new models. SDC [33] used the shift of current samples between two sessions $\delta_{i}^{t-1\to t}=e_{i}^{t}-e_{i}^{t-1}$ to estimate the update $\Delta_{c}^{t-1\rightarrow t}$. However, computing the offset between each training sample and all prototypes of old classes in the current stage is computationally expensive. As the number of classes increases, the computational cost will gradually escalate. A more efficient approach is to directly compute the offset between the prototypes on old model and the prototypes of the new model which can be represented as
$\varphi_{i}^{t-1\to t}=\varphi_{i}^{t}-\varphi_{i}^{t-1}$ (8)
We can compute $\varphi_{i}^{t-1}$ at the start of the current task with the model trained in task t1t-1. After training on the new task, we compute $\varphi_{i}^{t-1\to t}$ and use it to estimate $\Delta_{c}^{t-1\rightarrow t}$. So, we correct the semantic shift as
$\widetilde{\Delta}_{c}^{t-1\to t}=$ $\frac{\sum\alpha_{i}\varphi_{i}^{t-1\to t}}{\sum\alpha_{i}},c\notin C^{t},$ (9)
[…]
Where, the offset weight $\alpha$ is replaced by the cosine similarity between old and new class prototypes. The semantic shift is replaced by the offset of the new class prototype on the old and new model. Overall, before retraining the classifier, we update the prototypes with
[…]
where $N_{c}$ denotes the number of images in class $c$. The detailed computational steps are listed in algorithm 1.
@@ 4.3 Unified classifier training
$\mathcal{L}(\theta_{cls},\mathcal{V}_{c})=-\sum_{i=1}^{S_{n}*C}log\frac{\mathbf{e}^{(\theta_{cls}^{j}(v_{i}))}}{\sum_{k\in C}\mathbf{e}^{(\theta_{cls}^{k}(v_{i}))}},$ (11)
@@
Algorithm 2 Training Algorithm for SSIAT.
Input: The total number of sessions $T$; The pre-trained ViT model with classifier $M_{\theta}$; The sequence data $D$.
Output: The model $M_{\theta}$ after training in each learning session.
1: Initialize the classifier $f_{\theta_{cls}}$ of the model $M_{\theta}$
2:for t=1,,Tt=1,...,T do
3:    if $t=1$ then
4:     Train model $M_{\theta}$ using $\mathcal{L}^{t}$ with $D_{t}$ to get $M_{\theta}^{t}$
5:     Compute the Prototypes of each class using formula 7
6:    else
7:     Train model $M_{\theta}$ using $\mathcal{L}^{t}$ with $D_{t}$ to get $M_{\theta}^{t}$
8:     Compute the Prototypes of each new class using formula 7
9:     Update old class Prototypes with the algorithm of Semantic Shift Estimation.
10:     Construct Gaussian distribution for each class with updated Prototypes and then sample feature samples $s_{t}$
11:     Retrain the whole classifier $f_{\theta_{cls}}$ using formula 11 with feature samples $s_{t}$
12:end for
In general, our training process consists of two stages for each incremental task. First, we fine-tune the shared adapter in the ViT backbone to update feature expression. Then, we employ the classifier retraining which utilizes the Gaussian distribution sampling of feature representations for each class. This process involves optimizing the prototype offset of the old class and utilizing sampling from class Gaussian distributions. The specific algorithm is described in algorithm 2.
TABLE I: Experimental results on four CIL benchmarks. All other methods are reproduced using the same seeds for a fair comparison.
Method Params ImageNetR ImageNetA CUB200 CIFAR100
@@
RANPAC[76] 3.84M $77.83_{\pm 0.23}$ $82.97_{\pm 0.33}$ $60.46_{\pm 0.89}$ $68.11_{\pm 2.95}$ $89.95_{\pm 0.65}$ $93.67_{\pm 0.59}$ $91.58_{\pm 0.22}$ $94.44_{\pm 0.91}$
EASE [19] 11.9M $76.04_{\pm 0.16}$ $81.38_{\pm 0.36}$ $55.15_{\pm 0.78}$ $64.15_{\pm 1.47}$ $85.21_{\pm 0.50}$ $90.92_{\pm 0.76}$ $88.25_{\pm 0.54}$ $92.02_{\pm 1.04}$
SSIAT[34] 1.19M $79.38_{\pm 0.59}$ $83.63_{\pm 0.43}$ $62.43_{\pm 1.63}$ $70.83_{\pm 1.63}$ $88.75_{\pm 0.38}$ $93.00_{\pm 0.90}$ $91.35_{\pm 0.26}$ $94.35_{\pm 0.60}$
Ours 1.19M $79.61_{\pm 0.40}$ $83.71_{\pm 0.42}$ $62.73_{\pm 1.01}$ $71.05_{\pm 1.73}$ $88.32_{\pm 0.48}$ $92.95_{\pm 0.55}$ $91.51_{\pm 0.20}$ $94.29_{\pm 0.88}$
Figure 5: The performance of each learning session on six CIL datasets. (a) ImageNetR; (b) ImageNetA; (c) CUB200; (d) CIFAR100; (e) Stanford Cars; (f) Food-101 . These curves are plotted by calculating the average performance across three different seeds for each incremental session.
@@ 5.1 Datasets and Evaluation Protocols
Dataset: We evaluate our method on six commonly used CIL benchmarks and one cross-domain CIL dataset. We randomly split the dataset into 10 or 20 learning sessions in the CIL setting. Furthermore, we evaluate our method on few-shot class-incremental learning and hierarchical class-incremental learning datasets. CIFAR100 [77] is a widely used dataset in CIL which consists of 60000 images, belonging to 100 different categories. CUB200 [78] is a dataset that contains approximately 11,788 images of 200 bird species with fine-grained class labels. Additionally, we also follow recent work [15, 18] to use the other three datasets which have a large domain gap with pre-training data. ImageNetR [79] consists of 30,000 images with 200 categories. Although its categories overlap with ImageNet-21K [80], the images belong to a different domain. ImageNetA [81] is a real-world dataset that consists of 200 categories. This dataset exhibits significant class imbalance, with some categories having only a few training samples. Stanford Cars [82] is a fine-grained car dataset that consists of 196 categories including a total of 16185 images, and the training and test sets are divided in a 1:1 ratio. Food-101 [83] is a dataset containing 101 fine-grained foods with a total of 101k images. These two datasets are fine-grained datasets, which have a large domain gap with the pre-training data. VTAB [84] is a complex dataset that consists of 19 tasks covering a broad spectrum of domains and semantics. We follow previous work [15] to select 5 tasks to construct a cross-domain CIL dataset.
TABLE II: Experimental results for long-sequences (20 incremental sessions) on ImageNetR and ImageNetA dataset.
@@
EASE [19] $73.82_{\pm 0.55}$ $80.31_{\pm 0.70}$ $44.41_{\pm 2.65}$ $56.23_{\pm 3.46}$
SSIAT (Ours) $75.67_{\pm 0.14}$ $82.30_{\pm 0.36}$ $58.92_{\pm 1.82}$ $67.73_{\pm 0.25}$
TABLE III: Experimental results for fine-grained dataset on the CIL setting.
Method Food101 StanfordCars
SLCA [18] $86.66_{\pm 0.22}$ $91.57_{\pm 0.36}$ $49.58_{\pm 1.93}$ $62.32_{\pm 1.37}$
Adam-adapter[15] $3.51_{\pm 0.25} 88$$8.86_{\pm 0.42}$ $38.60_{\pm 0.15}$ $50.66_{\pm 0.59}$
Adam-ssf[15] $80.60_{\pm 0.32}$ $86.79_{\pm 0.40}$ $44.96_{\pm 0.62}$ $56.94_{\pm 0.65}$
Adam-prompt[15] $2.44_{\pm 0.92} 88$$8.14_{\pm 0.25}$ $37.37_{\pm 5.95}$ $50.47_{\pm 5.58}$
LAE [59] $82.26_{\pm 0.49}$ $87.88_{\pm 1.94}$ $41.87_{\pm 1.55}$ $57.78_{\pm 1.50}$
L2P [16] $79.40_{\pm 1.41}$ $85.70_{\pm 0.86}$ $32.87_{\pm 1.09}$ $48.02_{\pm 2.25}$
DualPrompt [17] $78.13_{\pm 0.12}$ $84.86_{\pm 0.96}$ $32.87_{\pm 1.09}$ $48.02_{\pm 2.25}$
CODAPrompt [28] $3.21_{\pm 0.10} 88$$8.85_{\pm 0.40}$ $39.01_{\pm 3.27}$ $53.71_{\pm 1.79}$
EASE [19] $81.99_{\pm 0.32}$ $87.95_{\pm 0.61}$ $36.46_{\pm 1.02}$ $62.32_{\pm 1.37}$
PANPAC [19] $89.02_{\pm 0.99}$ $92.78_{\pm 0.23}$ $2.79_{\pm 39.54} 55$$5.54_{\pm 39.17}$
SSIAT (Ours) $89.38_{\pm 0.11}$ $93.22_{\pm 0.40}$ $67.43_{\pm 1.47}$ $75.53_{\pm 1.26}$
Implementation details: We use ViT-B/16 [13] as the pre-trained model, which is pre-trained on ImageNet-21K [80]. The initial learning rate is set as 0.01 and we use the cosine Anneal scheduler. In our experiments, we train the first session for 20 epochs and 10 epochs for later sessions. Following previous papers [15, 18], we use common evaluation metrics in CIL. Specifically, we report the last session accuracy $\mathcal{A}_{Last}$ and average accuracy of the whole incremental sessions $\mathcal{A}_{Avg}=\frac{1}{T}\sum_{i=1}^{T}\mathcal{A}_{i}$. We utilize three different seeds to generate three different class orders for evaluating various methods. We report the mean and standard deviation based on the three experiments. See codes [^1] https://github.com/HAIV-Lab/SSIAT for more details.
@@ 5.2 Experiment Results
For a fair comparison, we compare our methods with SOTA CIL methods based on the pre-trained vision transformer model. We compare our methods with prompt-based methods L2P [27], DualPrompt [17], CODAPrompt [28], fine-tuning methods SLCA [18], and adapter-based method [15, 59, 75, 19]. Joint training and Fine-tuning are regarded as lower and upper bound in the CIL setting. Tab. I shows $\mathcal{A}_{Avg}$ and $\mathcal{A}_{Last}$ with three different seeds on four CIL benchmarks.
@@
TABLE IV: Experimental results for different methods on VTAB dataset which contain 5 datasets from different domains.
@@
EASE [19] 96.00 92.43 92.66 92.76 93.43 93.46
SSIAT (Ours) 96.40 93.14 93.96 93.86 94.38 94.34
TABLE V: The experiment results of hierarchical class-incremental learning on the CIFAR dataset.
[…]
Adam-vpt-shallow [15] $\rightarrow 88.92$ 75.00 81.96
[…]
Figure 6: Experimental results of few-shot incremental learning on cifar100 and cub200 datasets.
TABLE VI: Experimental results for baselines with different efficient tuning methods on CIFAR100. We report the overall performance of each session and the average performance.
@@
ImageNetR &\& ImageNetA: We report the performance on ImageNetR and ImageNetA in Tab. I. These two datasets are more difficult due to the domain gap with the pre-training data. It can be seen that the performance of each method on these two datasets is lower than CIFAR100 and CUB200. Besides, we can see that SLCA outperforms other previous methods significantly on these two datasets. Notably, SLCA achieves an impressive last accuracy on ImageNetR, surpassing the other methods. In contrast, our method achieves SOTA-level performance on both datasets with fewer tuning parameters. Based on Fig. 5, the performance of our method is slightly higher than SLCA in several learning sessions with fewer tuning parameters on the ImageNetR dataset. On the ImageNetA dataset, our method achieves the last accuracy of 62.43%, surpassing SLCA by 1.39%. The average accuracy across all sessions is 70.83%, showing a 2% improvement.
Additionally, we evaluate the performance of each method under the condition of long sequences. In this setting, each session consists of only 10 classes, and the results are summarized in Tab. II. Our method also maintains excellent performance in terms of $\mathcal{A}_{Last}$ and $\mathcal{A}_{Avg}$. The performance of SLCA is highly dependent on the class order in which the training data appears, resulting in a substantial variance in $\mathcal{A}_{Last}$ on ImageNetA. In contrast, the Adam-based methods remain relatively stable in long-sequence settings. For Adam-SSF, the long sequence only leads to a nearly $2\%$ performance drop in ImageNetR. However, for SLCA, its performance drops by $5\%$ on ImageNetR and nearly $10\%$ on ImageNetA. In comparison, our method demonstrates excellent stability on long sequences and outperforms other methods by a large margin.
Food101 &\& StanfordCars: We report the results of Food101 and StanfordCars on Tab. III. These two datasets are fine-grained datasets, and it is challenging for models to achieve good recognition performance solely based on pre-trained knowledge on such datasets. It can be observed that both first session adaptation-based methods and prompt-based methods perform poorly on these two datasets. In contrast, methods retraining classifiers, such as SLCA and our proposed method, demonstrate significant improvements in recognition performance compared to these methods. Besides, our method addresses the semantic shift of class prototypes, resulting in improved performance on both datasets.
Comparison to traditional CIL methods: We conduct evaluations by comparing our approach to SOTA traditional CIL methods shown in Tab. X. We replace the Resnet backbone with the pre-trained ViT model for fair comparison. The results indicate that the performance of iCaRL tends to be inferior compared to SOTA model expansion methods and our proposed method, even when past samples are stored. It can be observed that methods such as Foster and Der, which dynamically expand feature extraction networks, achieve impressive results on ImageNetR. The average accuracy of these methods is only $2\%$ lower than our method. However, on ImageNetA, where there are few-shot samples for many classes, these methods exhibit low performance.
TABLE VII: Ablation results for unified classifier training and semantic shift estimation on ImageNetA. We report the overall performance of each session and the average performance. We run the experiments with three seeds and reported the average performance.
@@
w/ SSCA 82.66 80.84 72.94 71.26 68.97 66.78 63.69 63.64 63.12 61.62 69.55
TABLE VIII: Linear probing results of different training ways on four datasets. We retrain the classifier using all the data on the fixed-trained backbone.
Method CIFAR100 ImageNetR ImageNetA CUB200
@@
TABLE IX: Experimental results of different adapter structures. We report the average performance and standard deviation.
@@
Figure 7: The comparison of the time costs between prototype-based semantic shift and sample-based semantic shift. The left shows the time consumed by different datasets and the right shows the time consumed by different sessions.
5.3 More CIL Settings and Results
Cross-Domain Class-Incremental Learning. VTAB is a cross-domain CIL dataset where each task provides training data from a different domain. Based on the results presented in Tab. IV, it can be observed that both SLCA and our method perform well in cross-domain CIL. Specifically, in the last incremental stage, our method achieves an accuracy that is $12\%$ higher than the Adam-based methods. Adam-based methods only perform fine-tuning in the first task and are not able to adapt well to subsequent tasks on the cross-domain dataset. Compared to Adam-based methods, EASE exhibits excellent performance while needing more model storage for expanded adapters.
Few-Shot Class-Incremental Learning. We report the performance of our method on CIFAR100 and CUB200 datasets in the few-shot incremental learning (FSCIL) setting, shown in Fig. 6. In the few-shot class-incremental learning setting, the model in the first stage learns a significantly larger number of classes compared to the incremental sessions that follow, and in the subsequent sessions, each class is provided with only $5$ training samples. We run the first session adaptation-based methods and some SOTA methods in few-shot incremental learning. It can be seen that the methods using the pre-trained ViT network have significantly improved compared to other methods. Furthermore, our method shows clear advantages throughout the entire incremental process.
Hierarchical Class-Incremental Learning. We report the performance of our method on the CIFAR100 dataset in hierarchical class-incremental learning (HCIL) setting, shown in Fig. V. In contrast to the traditional setting of treating all classes equally in the incremental process, HIL requires that any new classes learned in new sessions are a subclass from the last session. After the entire incremental process, the classes learned form a hierarchical tree structure. The model need to provide correct classification at each level of the hierarchy. It can be observed that compared to the first session adaptation method and traditional incremental learning approaches, our method achieves higher performance at all levels of the hierarchical structure.